OpenStudy (anonymous):

What is the slope of the line that passes through the pair points (2, 5) and (8, 3)? 1/3 -1/3 3 -3 What is the slope of the line that passes through the pair points (-5.2, 8.7) and (-3.2, 2.7)? 2 3 -2 -3 What is the slope of the line that passes through the pair points (3/2, -2) and (-3, 7/3) -27/26 -26/17 26/27 27/26

hartnn (hartnn):

The slope of the line through points (x1,y1) and (x2,y2) is given by : \(\huge m=\frac{y_1-y_2}{x_1-x_2}\) now,just put the values and find m in each case.

OpenStudy (anonymous):

isn't it \[m=\frac{ y _{2}-y _{1} }{ x _{2}-x _{1} }\]??

OpenStudy (anonymous):

y2-y1 ----- x2-x1 Data is right

hartnn (hartnn):

they both are same thing you'll get the same answer

OpenStudy (anonymous):

oh ok...

OpenStudy (anonymous):

So for number 1 it would be.. 5^1 - 3^2 __________ 2^1 - 8^2

hartnn (hartnn):

why did you square them ? it'll be just \(\large \dfrac{5-3}{2-8}=...?\)

OpenStudy (anonymous):

Oh I got confused, I thought those were exponents.

hartnn (hartnn):

i mean we take x1 = 2 and y1 = 5 those are just sub-scripts, not exponents :)

OpenStudy (anonymous):

Oh okay thanks, yeah sorry got confused xD

hartnn (hartnn):

so what slope did u get ? m=..

OpenStudy (anonymous):

Let me work it out, hold on.

OpenStudy (anonymous):

-1/3 ...?

hartnn (hartnn):

absolutely correct! :)

OpenStudy (anonymous):

Thank you! I never really got slopes, this helped me. Again, thanks c:

hartnn (hartnn):

welcome ^_^

OpenStudy (anonymous):

Wait sorry but how do I find out this one? What is the slope of the line that passes through the pair points (3/2, -2) and (-3, 7/3)? -27/26 -26/17 26/27 27/26 Because I got 4.5 _____ -4.33

hartnn (hartnn):

[-2-7/3]/[3/2-(-3)]

hartnn (hartnn):

[-13/3]/ [9/2] = -13/3 times 2/9 simplify

OpenStudy (anonymous):

-26/27?

hartnn (hartnn):

yep :)

OpenStudy (anonymous):

Thanks again! cx

OpenStudy (anonymous):

Wish I could give you another metal! XD

hartnn (hartnn):

i am not here for medals.... if you understand, i am happy, thats my medal :)

OpenStudy (anonymous):

Haha, well that's great ^-^ Happiness is the biggest reward someone can get :D